At the café the total price of four fajita dinners and three burrito dinners is $48. and the total price of three fajita dinners and two burrito dinners is $34. What is the price of each type of dinner?
Answer by stanbon(75887)

At the café the total price of four fajita dinners and three burrito dinners is $48. and the total price of three fajita dinners and two burrito dinners is $34. What is the price of each type of dinner?
-----------
Equations:
4f + 3b = 48 dollars
3f + 2b = 34 dollars
-----------------------
Modify::
12f + 9b = 3*48
12f + 8b = 4*34
-------------------------
Subtract and solve for "b"::
b = 8 (price of one burrito dinner)
----
Solve for "f" using 3f + 2b = 34
3f + 16 = 34
3f = 18
f = 6 (price of one fajita dinner)
================
Cheers,
Stan H.
----------------